OpenStudy (migitmack):

Surf's Up charges $22 per hour to rent a surfboard and $8 per hour to rent a wetsuit. Marty and Cheryl paid a total of $240 to rent a surfboard and a wetsuit each for a few hours one day. Cheryl rented the surfboard and wetsuit for 2 hours less than Marty. The following equation represents this situation, where x represents the number of hours Marty rented the surfboard and the wetsuit. 22x + 8x + 22(x − 2) + 8(x − 2) = 240 What is the total amount Marty paid to rent the surfboard and wetsuit that day? $30 $90 $120 $150

OpenStudy (migitmack):

HELP PLEASD

OpenStudy (anonymous):

ok first of all solve the above equation to get no of hours marry rented. 22x+8x+22x-44+8x-16=240 60x=300 x=300/60 x=5 hrs

OpenStudy (anonymous):

now since you got no of hours. you know the prices per hour just Multiply them and there you are ! if you are unable to do this i will be back in 20 min.

OpenStudy (migitmack):

So i multiply? what

OpenStudy (anonymous):

I don't know, I'm sorry =( I'm half asleep....

OpenStudy (anonymous):

thats the sweet stuff right there ight.....imma hooka u up

OpenStudy (anonymous):

theres more too ^

OpenStudy (anonymous):

@Migitmack you here ??

OpenStudy (migitmack):

YES

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

medal plz

OpenStudy (anonymous):

are you ok with this that have found above . x=5 hrs is there any problem with this ?

OpenStudy (migitmack):

Yahoo has redeemed themselves

OpenStudy (anonymous):

@Migitmack x=5 hrs ???? is it ok.

OpenStudy (migitmack):

i got the answer its 150

OpenStudy (anonymous):

yes you are right !!!!!!!!!!

OpenStudy (anonymous):

Because 22*5+8*5=150